OpenStudy (anonymous):

Add and simplify a+2/3+a+2/4

OpenStudy (mertsj):

\[\frac{a+2}{3}\times\frac{4}{4}+\frac{a+2}{4}\times\frac{3}{3}=\frac{4a+8}{12}+\frac{3a+6}{12}\]

OpenStudy (mertsj):

Now the denominators are the same so add the numerators and keep the same denominator.